OpenStudy (gabbyalicorn):

The lid for a circular plastic food container has a diameter of 9 cm. What is the approximate area of the top of the lid? Use 3.14 to approximate pi and round your answer to the nearest hundredth. 2 x pi x 9?

OpenStudy (michele_laino):

the requested area is given by the subsequent formula: \[A = \pi \times \frac{{{9^2}}}{4} = 3.14 \times \frac{{81}}{4} = ...c{m^2}\]

OpenStudy (gabbyalicorn):

62.775 = 1,017.36

OpenStudy (michele_laino):

no, it is: \[ = \pi \times \frac{{{9^2}}}{4} = 3.14 \times \frac{{81}}{4} = 63.585\;c{m^2}\]

OpenStudy (gabbyalicorn):

63.585= 63.585 yeppers

OpenStudy (michele_laino):

now you have to round off to the nearest hundredth, what do you get?

OpenStudy (gabbyalicorn):

64.

OpenStudy (gabbyalicorn):

am i right?

OpenStudy (michele_laino):

I think not, for example if I have to round off this number to the nearest hundredth 1.156 then I get: 1.16

OpenStudy (michele_laino):

so, what do you get?

OpenStudy (gabbyalicorn):

oh 63.60

OpenStudy (michele_laino):

better is 63.59

OpenStudy (gabbyalicorn):

:D

OpenStudy (michele_laino):

:)

OpenStudy (gabbyalicorn):

Thankies!

OpenStudy (michele_laino):

:)